OpenStudy (anonymous):

The sum of two numbers is 57. One of the numbers is 21. Which of the following equations could be used to find the other number?

OpenStudy (anonymous):

x = 57-21

OpenStudy (anonymous):

so x = 36

hero (hero):

Yup

OpenStudy (anonymous):

it has these options n = 57 + 21 21 = n – 57 57 = 2 + 21 Not Here 57 = n + 21

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

57 = n + 21

hero (hero):

I agree

OpenStudy (anonymous):

thank u i would have never guessed that thanks

hero (hero):

QED, make a mistake. It'll help me feel better

OpenStudy (anonymous):

lol plz dont though

OpenStudy (anonymous):

:D